# Depth-Weighted Detection of Behaviours of Risk in People with Dementia using Cameras

### v1

@@ 3.4 Data Preprocessing
3.5 Behaviours of Risk Detection
Spatio-temporal convolutional autoencoders (3DCAE) learn to reconstruct the input videos by minimizing the reconstruction error during training. The intuition is that 3DCAE is trained to reconstruct only normal instances. During testing, normal instances will have a lower reconstruction error, while instances with high reconstruction error are treated as anomalies (behaviours of risk in our case). Previously, a customized 3DCAE was used to detect behaviours of risk from data from one participant and a single camera [11, 4]. However, previous approaches showed high false positives due to obstruction caused by activities such as crowded scenes, large objects, or other events happening near the camera due to reasons described in Section 1. Such instances are extremely common in normal day-to-day activities happening in care units. As such, there was a need to handle this issue to make the method suitable for real-world deployment. Hence in this work, we propose a novel depth-weighted loss function to train a convolutional autoencoder, depCAE, where the depth of the pixels is used as a weight factor to enforce equivalent importance to both behaviours of risk and non-behaviours of risk events happening far and near the camera. It is hypothesized that the weighted-depth loss function will lead to a reduction in false positives in this problem. We also propose a new method to determine the threshold for detecting behaviours of risk as anomalies using the abnormal non-behaviours of risk activities in the training set as outliers.
The depCAE followed an encoder-decoder architecture, where the encoder consisted of $2$D convolution and max-pooling layers, with a kernel size of ($1\times 3\times 3$), stride ($1\times 1\times 1$) and padding ($0\times 1\times 1$), followed by batch normalization and ReLU operations. The decoder used $2$D deconvolution layers ($2$D transposed convolution operation) with a kernel size of ($1\times 3\times 3$), strides ($1\times 1\times 1$), ($1\times 2\times 2$), ($1\times 2\times 2$) and paddings ($0\times 1\times 1$), ($0\times 1\times 1$), ($0\times 1\times 1$) for first, second, and third $2$D deconvolution layers, respectively. The overall architecture of the depCAE is presented in Fig. 2.
Depth-weighted Anomaly Score: Existing CAE methods [4] employ reconstruction error to detect anomalous frames. This strategy inherently favours anomalies closer to the camera, as they occupy more pixels in the image plane. This bias poses a challenge, especially in scenes with highly varying distances from the camera, such as corridors (see Fig. 1). As a result, anomalies at different distances receive unequal anomaly scores, with those closer to the camera being unfairly favoured. To address this problem, we propose a loss function where depth per pixel is used as weight during the calculation of reconstruction error to enforce equivalent importance to both near and far pixels.
[…]
Figure 3: (a) Illustration of the geometric relationship between volume and area of the projected square into the image plane. (b) Derivation of the relationship between projected areas at different distances from the camera.
To understand the translation of an anomaly in $3$D space, we describe the relationship between the area $A$ depicting the anomaly in the image and its $3$D volume. For simplicity, consider this volume as a square prism with dimensions equal to $L$ (see Fig. 3 (a)). Consider now two of the vertex of the square prism represented by two $3$D points $x_{1}=(X_{1},Y_{1},Z_{1})$ and $x_{2}=(X_{2},Y_{2},Z_{2})$ in the $3$D space and their $2$D projections $u_{1}=(U_{1},V_{1})$ and $u_{2}=(U_{2},V_{2})$. The projection of a $3$D point onto a $2$D image plane is expressed by the image coordinates $(U,V)$, related to the $3$D world coordinates $(X,Y,Z)$ through the pinhole camera model equations:
$U=\frac{fX}{Z},\quad V=\frac{fY}{Z}$ (1)
where $f$ is the focal length and $Z$ is the distance from the $3$D point to the camera position. Assuming the difference in depth $Z$ between these points is negligible ($Z\approx Z_{1}\approx Z_{2}$), the projections on the image plane according to equation (1) are:
$U_{1}=\frac{fX_{1}}{Z},\quad U_{2}=\frac{fX_{2}}{Z}$ (2)
Given that the distance between the points in the 3D space along the X-axis is $L=|X_{2}-X_{1}|$, the area $A$ of the square in the $2$D plane containing the anomaly is:
To analyze the sensitivity of the anomaly score $(AS)$ on the distance from the camera and therefore, on the area in the image plane, we can express the anomaly score $(AS)^{A}$ of an area $A$ as:
$(AS)^{A}=\overline{(AS)^{\text{pixel}}}\cdot A$ (4)
where $\overline{(AS)^{\text{pixel}}}$ represents the average anomaly score per pixel, a constant value that depends solely on the anomaly itself, irrespective of its position in 3D space or its projection onto the image plane.
Consider the same anomaly event, i.e., same value of $\overline{(AS)^{\text{pixel}}}$, at different distances $Z_{1}$ and $Z_{2}$ from the camera. As illustrated in Fig. 3 (b), this results in different area values $A_{1}$ and $A_{2}$ in the image plane, with $A_{1}<A_{2}$ if $Z_{1}~{}>~{}Z_{2}$. Since both anomalies have the same value of $\overline{(AS)^{\text{pixel}}}$, according to equation (4), the difference in area will result in a different anomaly score. To ensure depth invariance, we need the anomaly to have the same score $AS_{inv}$ regardless of its distance from the camera. This is achieved by introducing a factor $K$:
$AS_{inv}=K_{1}\cdot(AS)^{A_{1}}=K_{2}\cdot(AS)^{A_{2}}$ (5)
Combining the equations (3), (4) and (5), and since $\overline{(AS)^{\text{pixel}}}$ is a constant, we can express $AS_{inv}$ as :
$AS_{inv}=K_{1}\left(fL\right)^{2}\frac{1}{Z_{1}^{2}}=K_{2}\left(fL\right)^{2}\frac{1}{Z_{2}^{2}}$ (6)
Therefore, given the constant value of $\left(fL\right)^{2}$, to achieve depth invariance, the factors $K_{1}$ and $K_{2}$ must satisfy the following conditions:
$K_{1}=Z_{1}^{2},\quad K_{2}=Z_{2}^{2}$ (7)
Based on this, we propose the following novel loss function where depth per pixel is used as weight during the calculation of reconstruction error as a depth-weighted mean squared error to enable depCAE to enforce equivalent importance to both near and far pixels.
$L_{dmse}=\frac{1}{N_{e}}\sum_{l=1}^{W}\sum_{i=1}^{S}\sum_{j=1}^{S}Z_{l,i,j}(I_{l,i,j}-O_{l,i,j})^{2}$ (8)
where, $I$ is the input pixel, $O$ is the reconstructed pixel, $Z$ is the pixel depth, $S$ is the spatial size, $W$ represents the number of frames in a window and is termed as the window size and $N_{e}$ is the total number of pixels in a window. In the depCAE method, $W$ = 75 and $N_{e}$ = $75\times 64\times 64=307200$. Empirically, we found that using a weight $Z$ in equation (8) gave better performance than the weight $Z^{2}$. The weighted-depth loss function helped the method to treat both the near and far behaviours of risk events with similar importance, which in turn helped to reduce false positives. The depth map was estimated using the work by Ranftl et al. [19]. During testing, the depth-weighted mean squared error was used as anomaly score for detecting behaviours of risk.
Threshold Determination: In an anomaly detection approach, it can be challenging to determine an operating threshold due to the absence of anomalous samples in the training set. As discussed in Section 3.3, the data was annotated for other unusual activities, such as large objects, crowded scenes and other infrequent activities in the training set. Due to the unusual nature of these annotated activities relative to the activities of daily living in a dementia unit, we propose to treat these activities as proxy outliers to create a proxy validation set to determine the threshold. Following a similar approach as Khan et al. [20], we chose the training reconstruction error with maximum F1-score on proxy validation set as the operating threshold.
Performance Evaluation: The primary metric employed for evaluating the performance of methods was the area under the curve of receiver operating characteristic (AUROC), supplemented by F1-score. Additional metrics are also reported, including true positive rate (TPR), true negative rate (TNR), false positive rate (FPR), false negative rate (FNR), gmean ($\sqrt{TPR*TNR}$), Matthew’s Correlation Coefficient (MCC), precision, recall, specificity, and balanced accuracy [21]. Transformers and variational autoencoders have performed well in the existing literature for the task of VAD [22, 23]. Hence, in addition to the baseline 3DCAE method [4], the performance of the depCAE for detecting behaviours of risk in PwD as anomalies was compared with two additional existing methods, namely a self-supervised masked convolutional transformer [22] block-based autoencoder 3DCAE_SSMCTB and a variational autoencoder VAE [24]. The 3DCAE_SSMCTB method used a self-supervised masked convolutional transformer block [22] embedded after the encoder layer and its output was passed as input to the decoder. The transformer block [22] was structured as a convolutional layer with dilated masked kernels, followed by a transformer module that performed channel attention. A self-supervised loss function was used with the block to minimize the reconstruction error between the masked input and the predicted output. VAE [24] learned to generate two vectors that represent the parameters (mean and variance) of a distribution from which the latent vector was sampled instead of learning how to generate a latent vector that the decoder function can reproduce. For fairness during comparison, the encoder and decoder architecture of 3DCAE, 3DCAE_SSMCTB and VAE was kept the same as depCAE and they used interquartile range (IQR) generated proxy-validation set [20] for threshold calculation.
Table 4: Performance of different methods for behaviours of risk detection in PwD on three different cameras.
Metric Cam1 Cam2 Cam3
3DCAE 3DCAE_ SSMCTB VAE depCAE 3DCAE 3DCAE_ SSMCTB VAE depCAE 3DCAE 3DCAE_ SSMCTB VAE depCAE
TPR 0.945 0.950 0.641 0.939 0.949 0.728 0.502 0.844 0.919 0.649 0.230 0.730
TNR 0.628 0.581 0.713 0.684 0.611 0.668 0.812 0.686 0.530 0.645 0.903 0.714
FPR 0.372 0.419 0.287 0.316 0.389 0.332 0.188 0.314 0.470 0.355 0.097 0.286
FNR 0.055 0.050 0.359 0.061 0.051 0.272 0.498 0.156 0.081 0.351 0.770 0.270
Gmean 0.770 0.743 0.676 0.801 0.761 0.697 0.638 0.761 0.698 0.647 0.455 0.722
MCC 0.301 0.276 0.198 0.335 0.295 0.215 0.200 0.290 0.220 0.149 0.105 0.235
Precision 0.165 0.150 0.148 0.187 0.162 0.148 0.174 0.176 0.118 0.111 0.139 0.149
Recall 0.945 0.950 0.641 0.939 0.949 0.728 0.502 0.844 0.919 0.649 0.230 0.730
Specificity 0.628 0.581 0.713 0.684 0.611 0.668 0.812 0.686 0.530 0.645 0.903 0.714
Balanced Accuracy 0.786 0.765 0.677 0.811 0.780 0.698 0.657 0.765 0.724 0.647 0.566 0.722
F1-score 0.281 0.259 0.240 0.313 0.276 0.246 0.259 0.291 0.210 0.190 0.173 0.248
AUROC 0.844 0.779 0.754 0.852 0.810 0.766 0.790 0.810 0.765 0.739 0.709 0.768
@@ 3.5 Behaviours of Risk Detection
Table 5: Ablation analysis of proposed components across three different cameras.
@@
4 Results and Discussion
In this section, we present the results to gauge the effectiveness of depCAE in detecting behaviours of risk in PwD and compare its performance with existing methods 3DCAE, 3DCAE_SSMCTB and VAE. We further conducted an ablation analysis of different components of the depCAE method and performed participant and sex-specific analysis of depCAE performance.
Performance Comparison: Table 5 shows the performance of depCAE compared to other methods for the detection of behaviours of risk in PwD across three cameras. The best values for the primary performance metrics AUROC and F1-score are marked in bold. The cases where depCAE gave lower FPR as compared to baseline 3DCAE method are marked in gray cells. Our depCAE method achieved the best AUROC and F1-score for all three cameras in comparison to other methods. The depCAE method achieved an AUROC and F1-score of $0.852$ and $0.313$ for Cam1, $0.81$ and $0.291$ for Cam2, and $0.768$ and $0.248$ for Cam3. Further, our depCAE method gave lower FPR as compared to 3DCAE and 3DCAE_SSMCTB methods for all three cameras. This shows that the proposed idea of using pixel depth as weight in the loss function is indeed effective in reducing false positives. The depCAE method performed better for Cam1 and Cam2 in comparison to Cam3, which can be attributed to the major portion of Cam3 view being occluded by the wall. This made it difficult to gauge the behaviours of risk events clearly in Cam3.
Ablation Analysis: We investigated the relative effectiveness of the two components of the proposed method, namely, depth-weighted loss function and threshold determination using annotated outliers, in detecting behaviours of risk in PwD. Table 5 presents the results for ablation analysis over three cameras, where depWgtOnly refers to the setting of depCAE trained using depth-weighted loss function while using IQR outliers [20] for threshold calculation, and anntThrOnly refers to the setting of using mean squared error loss function while using annotated outliers for threshold calculation. The best AUROC and F1 scores are marked in bold. For Cam1, anntThrOnly performed better than depWgtOnly in terms of F1-score, while depWgtOnly performed better in terms of AUROC. Overall, depCAE benefited from combined performance of depWgtOnly and anntThrOnly giving an AUROC of $0.852$. For Cam2, anntThrOnly performed similar to depCAE in terms of AUROC, while the performance of depCAE degraded in comparison to anntThrOnly in terms of F1-score. For Cam3, depCAE benefited from the combined functionality of depWgtOnly and anntThrOnly, outperforming both of them in terms of F1-score and AUROC. Overall, it can be observed that both depWgtOnly and anntThrOnly are vital components of depCAE and contributed significantly towards the performance of depCAE.
@@ 4 Results and Discussion
Participant-Specific Analysis: As presented in Table 3, the type of behaviours of risk varied across participants. Hence, it is crucial to consider the potential dissimilar performance across individuals when training and evaluating behaviours of risk detection methods. Table 7 presents the performance of depCAE separately for each participant involved in the study across three cameras. “-” represents the cases where one or less than one agitation window was recorded for the participants. In general, it can be observed that depCAE showed better behaviours of risk detection performance for participants showing less diverse episodes of behaviours of risk, namely, participants $2$, $7$ and $8$, as compared to participants with more diverse episodes. depCAE performed generally well for participant $3$ across cameras $1$ and $2$ and across all three cameras for participant $4$, for whom the most number of behaviours of risk windows were recorded. On average, the detection performance was better for participants on cameras $1$ and $2$ as compared to camera $3$.
@@
Sex-Specific Analysis: It is important to investigate the performance of depCAE across population groups. Table 7 presents the performance of depCAE stratified based on sex across three cameras. A similar behaviour of risk detection performance was observed for both males and females for Cam1. However, for cameras $2$ and $3$ it was observed that the depCAE method performed better for females than males. The reason for this can be attributed to the presence of more females than males in the study, leading to the skewed class distribution due to the presence of more behaviours of risk windows for females than males, enhancing the model’s learning and generalization capabilities for females.

### v2

@@ 3.5 Behaviours of Risk Detection
Experimental Setup: Transformers and variational autoencoders have performed well in the existing literature for VAD [27, 28]. In this work, we have used three different anomaly detection methods, namely, 3DCAE, a self-supervised masked convolutional transformer block-based autoencoder (CTBAE) and a variational autoencoder (VAE) for evaluation. The 3DCAE method has been adapted from our previous work [4]. The CTBAE method used a self-supervised masked convolutional transformer block [27] embedded after the encoder layer and its output was passed as input to the decoder. VAE [29] learned to generate two vectors that represent the parameters (mean and variance) of a distribution from which the latent vector was sampled. For fairness during comparison, the encoder and decoder architecture of 3DCAE, CTBAE and VAE were kept the same and they used interquartile range (IQR) generated proxy-validation set [26] for threshold calculation. The proposed depth-weighted approach was integrated as part of the 3DCAE, CTBAE and VAE methods and their depth variants are referred to as depCAE, depCTBAE and depVAE, respectively. The depCAE method followed an encoder-decoder architecture, where the encoder consisted of $2$D convolution and max-pooling layers, with a kernel size of ($1\times 3\times 3$), stride ($1\times 1\times 1$) and padding ($0\times 1\times 1$), followed by batch normalization and ReLU operations. The decoder used $2$D deconvolution layers ($2$D transposed convolution operation) with a kernel size of ($1\times 3\times 3$), strides ($1\times 1\times 1$), ($1\times 2\times 2$), ($1\times 2\times 2$) and paddings ($0\times 1\times 1$), ($0\times 1\times 1$), ($0\times 1\times 1$) for first, second, and third $2$D deconvolution layers, respectively. The overall architecture of the depCAE is presented in Fig. 3. In the depCAE method, $W$ = 75 and $N_{e}$ = $75\times 64\times 64=307200$.
Evaluation Metrics: The proposed approach aims to reduce false positives; however, it is also important to consider false negatives when detecting behaviours of risk in PwD. Hence, when evaluating performance, it is important to consider a metric that can give a balanced estimate of the performance on both false positive rate (FPR) and false negative rate (FNR). Precision and recall can be expressed as,
$Precision=\frac{TPR}{TPR+\frac{1}{\gamma}FPR}$ (9)
$Recall=\frac{TPR}{TPR+FNR}$ (10)
where TPR refers to true positive rate and $\gamma=\frac{No.\>of\>positives\>(P)}{No.\>of\>negatives\>(N)}$ is class ratio. As seen in eqns. 9 and 10, precision and recall give an estimate of the performance over FPR and FNR, respectively. F1-score is the harmonic mean of precision and recall, and provides a reasonable performance estimate over TPR, FPR and FNR, simultaneously. Matthew’s Correlation Coefficient (MCC) gives an overall performance estimate over true positives (TP), true negatives (TN), false positives (FP) and false negatives (FN) and is expressed as,
$MCC=\frac{TP*TN-FP*FN}{\sqrt{(TP+FP)*P*N*(TN+FN)}}$ (11)
Both F1-score and MCC are threshold-based metrics; hence they evaluate performance over a single threshold. The area under the curve of receiver operating characteristic (AUROC) evaluates the performance over TPR and FPR over a range of thresholds. Hence, the primary metrics employed for evaluating the performance of methods in this work are MCC, F1-score and AUROC. We also report additional metrics for reference, namely, TPR, true negative rate (TNR), FPR, FNR, gmean ($\sqrt{TPR*TNR}$), precision, recall, specificity, and balanced accuracy [30]. The evaluation metrics use reconstruction error as the score to identify behaviours of risk as anomalies during testing.
Table 4: Performance comparison of depth variants and existing methods for Cam1.
Existing methods Depth variants
3DCAE CTBAE VAE depCAE depCTBAE depVAE
TPR 0.945 0.950 0.641 0.939 0.773 0.552
TNR 0.628 0.581 0.713 0.684 0.742 0.780
FPR 0.372 0.419 0.287 0.316 0.258 0.220
FNR 0.055 0.050 0.359 0.061 0.227 0.448
Gmean 0.770 0.743 0.676 0.801 0.758 0.657
Precision 0.165 0.150 0.148 0.187 0.189 0.163
Recall 0.945 0.950 0.641 0.939 0.773 0.552
Specificity 0.628 0.581 0.713 0.684 0.742 0.780
Balanced Accuracy 0.786 0.765 0.677 0.811 0.758 0.666
MCC 0.301 0.276 0.198 0.335 0.293 0.200
F1-score 0.281 0.259 0.240 0.313 0.304 0.252
AUROC 0.844 0.779 0.754 0.852 0.823 0.784
Table 5: Performance comparison of depth variants and existing methods for Cam2.
Existing methods Depth variants
3DCAE CTBAE VAE depCAE depCTBAE depVAE
TPR 0.949 0.728 0.502 0.844 0.412 0.751
TNR 0.611 0.668 0.812 0.686 0.832 0.719
FPR 0.389 0.332 0.188 0.314 0.168 0.281
FNR 0.051 0.272 0.498 0.156 0.588 0.249
Gmean 0.761 0.697 0.638 0.761 0.586 0.735
Precision 0.162 0.148 0.174 0.176 0.163 0.175
Recall 0.949 0.728 0.502 0.844 0.412 0.751
Specificity 0.611 0.668 0.812 0.686 0.832 0.719
Balanced Accuracy 0.780 0.698 0.657 0.765 0.622 0.735
MCC 0.290 0.215 0.200 0.290 0.164 0.264
F1-score 0.276 0.246 0.259 0.291 0.234 0.284
AUROC 0.810 0.766 0.790 0.810 0.757 0.792
Table 6: Performance comparison of depth variants and existing methods for Cam3.
Existing methods Depth variants
3DCAE CTBAE VAE depCAE depCTBAE depVAE
TPR 0.919 0.649 0.230 0.730 0.743 0.135
TNR 0.530 0.645 0.903 0.714 0.622 0.931
FPR 0.470 0.355 0.097 0.286 0.378 0.069
FNR 0.081 0.351 0.770 0.270 0.257 0.865
Gmean 0.698 0.647 0.455 0.722 0.680 0.355
Precision 0.118 0.111 0.139 0.149 0.119 0.119
Recall 0.919 0.649 0.230 0.730 0.743 0.135
Specificity 0.530 0.645 0.903 0.714 0.622 0.931
Balanced Accuracy 0.724 0.647 0.566 0.722 0.682 0.533
MCC 0.220 0.149 0.105 0.235 0.182 0.063
F1-score 0.210 0.190 0.173 0.248 0.205 0.127
AUROC 0.765 0.739 0.709 0.768 0.722 0.696
4 Results and Discussion
In this section, we evaluate the performance of the proposed approach for FP reduction and detecting behaviours of risk in PwD. We conduct an ablation analysis of different components of the proposed approach and the effect of frame size and frame rate on the performance of detecting behaviours of risk. We further perform cross-camera, participant and sex-specific analysis.
Performance Evaluation: We investigate the effectiveness of the proposed approach for reducing FP and detecting behaviours of risk by evaluating the improvement in performance after integrating the proposed approach into existing methods. As discussed in Section 3.5, depth variants were created after integrating the proposed approach into the existing methods. Tables 6, 6 and 6 compare the performance of the depth variants with their existing method counterparts for behaviours of risk detection for Cam1, Cam2 and Cam3, respectively. The best values for FPR and primary evaluation metrics, MCC, F1-score and AUROC, are marked in bold. It can be observed that depCAE performed mostly better than 3DCAE in terms of MCC, F1-score and AUROC for all three cameras. depCTBAE performed better than CTBAE for Cam1 and Cam3, while depVAE performed better than VAE for Cam1 and Cam2. depCAE recorded lower FPR than 3DCAE for all three cameras. depCTBAE and depVAE recorded lower FPR than CTBAE and VAE, respectively, for two out of three cameras. The methods performed better for Cam1 and Cam2 in comparison to Cam3, which can be attributed to the major portion of Cam3 view being occluded by the wall. This made it difficult to gauge the behaviours of risk events clearly in Cam3. For the case of secondary evaluation metrics, gmean, precision and specificity, the depth variants mostly performed better than the existing methods across the three cameras. The depCAE method achieved the best MCC, F1-score and AUROC overall for all three cameras; hence further analysis in this section is performed using depCAE method. The depCAE method achieved a MCC, F1-score and AUROC of $0.335$, $0.313$, $0.852$ for Cam1, $0.29$, $0.291$, $0.81$ for Cam2, and $0.235$, $0.248$, $0.768$ for Cam3, respectively. While depVAE gave lower FPR as compared to other methods, it gave significantly higher FNR. In such cases, it is important to consider a balanced evaluation over both FPR and FNR, which is provided by F1-score, where depCAE performed the best. It can be inferred from the observations that the proposed approach of weighing the loss with pixel depth is indeed effective in reducing false positives and improving detection performance. Further, the proposed approach can be easily integrated into any existing autoencoder-based anomaly detection method.
Ablation Analysis I (Relative Effectiveness of the Proposed Components): We investigate the relative effectiveness of the two components of the proposed method, namely, depth-weighted loss and threshold determination using annotated outliers, in detecting behaviours of risk in PwD. Table 7 presents the results for ablation analysis over three cameras, where depWgtOnly refers to the setting of depCAE trained using depth-weighted loss while using IQR outliers [26] for threshold calculation, and anntThrOnly refers to the setting of using mean squared error loss while using annotated outliers for threshold calculation. The best MCC, F1 scores and AUROC are marked in bold. For Cam1, anntThrOnly performed better than depWgtOnly in terms of MCC and F1-score, while depWgtOnly performed better in terms of AUROC. Overall, depCAE benefited from the combined performance of depWgtOnly and anntThrOnly giving an AUROC of $0.852$. For Cam2, anntThrOnly performed similar to depCAE in terms of AUROC, while the performance of depCAE degraded in comparison to anntThrOnly in terms of MCC and F1-score. For Cam3, depCAE benefited from the combined functionality of depWgtOnly and anntThrOnly, outperforming both of them in terms of MCC, F1-score and AUROC. Overall, it can be observed that both depWgtOnly and anntThrOnly are vital components of depCAE and contribute significantly towards the performance of depCAE.
Ablation Analysis II (Effect of Frame Size and Frame Rate): We investigate the effect of increasing the frame size and decreasing the frame rate in detecting behaviours of risk in PwD. Table 8 presents the results for ablation analysis over three cameras, where s64r15 refers to the original $64\times 64$ frame size and $15$ fps setting, and s128r8 refers to $128\times 128$ frame size and $8$ fps setting. For s128r8, a lower frame rate was used to manage the computation complexity added due to increase in frame size. The best MCC, F1 score and AUROC are marked in bold. It was observed that for Cam1, s128r8 performed better than s64r15 in terms of MCC and F1-score, while s64r15 performed better in terms of AUROC. For Cam2 and Cam3, s64r15 performed better in terms of all of MCC, F1-score and AUROC metrics. Overall, it was observed that s64r15 performed similar or better than s128r8. Increasing the frame size did not lead to any significant improvement in performance. The depCAE method can extract the necessary information for detecting behaviours of risk from a low resolution of $64\times 64$, which may be difficult for the human eye.
@@ 4 Results and Discussion
Table 7: Ablation analysis of proposed components of depCAE across three different cameras.
@@
Table 8: Ablation analysis of input frame size and frame rates for depCAE.
Metric Cam1 Cam2 Cam3
s64r15 s128r8 s64r15 s128r8 s64r15 s128r8
Gmean 0.801 0.808 0.761 0.645 0.722 0.653
Precision 0.187 0.199 0.176 0.198 0.149 0.105
Recall 0.939 0.917 0.844 0.494 0.730 0.851
Specificity 0.684 0.712 0.686 0.841 0.714 0.501
Balanced Accuracy 0.811 0.815 0.765 0.668 0.722 0.676
MCC 0.335 0.345 0.290 0.226 0.235 0.173
F1-score 0.313 0.326 0.291 0.282 0.248 0.187
AUROC 0.852 0.848 0.810 0.795 0.768 0.739
Table 9: Cross-camera analysis for depCAE.
Train-test on Cam1 Train on Cam2 test on Cam1 Train-test on Cam2 Train on Cam1 test on Cam2
Gmean 0.801 0.750 0.761 0.704
Precision 0.187 0.162 0.176 0.139
Recall 0.939 0.862 0.844 0.848
Specificity 0.684 0.653 0.686 0.585
Balanced Accuracy 0.811 0.757 0.765 0.717
MCC 0.335 0.274 0.290 0.227
F1-score 0.313 0.272 0.291 0.239
AUROC 0.852 0.796 0.810 0.770
Cross-Camera Analysis: We investigate the performance of depCAE to detect behaviours of risk in a new camera that was not seen during training. As part of this analysis, the depCAE method was trained on Cam2 and tested on Cam1 and vice-versa, as seen in Table 9. The results are compared with the setting where the method is trained and tested on the same camera. It was observed that there was a drop in performance when the method was trained on a camera different from the testing camera. A drop in F1-score performance of $0.041$ ($0.313$-$0.272$) and $0.052$ ($0.291$-$0.239$), and in AUROC performance of $0.056$ ($0.852$-$0.796$) and $0.04$ ($0.81$-$0.77$) was observed when the method was tested on Cam1 and Cam2, respectively. The performance drop was expected and within reasonable limits as the training camera was different from the testing camera and no fine-tuning was involved. We have discussed some strategies to improve the cross-camera performance in Section 5 as part of our future work.
@@
Participant-Specific Analysis: As presented in Table 3, the type of behaviours of risk varied across participants. Hence, it is crucial to consider the potential dissimilar performance across individuals when training and testing behaviours of risk detection methods. Table 11 presents the performance of depCAE separately for each participant involved in the study across three cameras. “-” represents the cases where one or less than one agitation window was recorded for the participants. In general, it can be observed that depCAE showed better behaviours of risk detection performance for participants showing less diverse episodes of behaviours of risk, namely, participants $2$, $7$ and $8$, as compared to participants with more diverse episodes. depCAE performed generally well for participant $3$ across cameras $1$ and $2$ and across all three cameras for participant $4$, for whom the most number of behaviours of risk windows were recorded. On average, the detection performance was better for participants on cameras $1$ and $2$ as compared to camera $3$.
@@
Sex-Specific Analysis: It is important to investigate the performance of depCAE across population groups. Table 11 presents the performance of depCAE stratified based on sex across three cameras. As can be seen in Table 3, some participants did not show any behaviours of risk event in one or two cameras, leading to the situation with only 6 participants with differing participant ids for any individual camera. However, the overall sex-specific analysis was done for all 9 participants. A similar behaviour of risk detection performance was observed for both males and females for Cam1. However, for cameras $2$ and $3$ it was observed that the depCAE method performed better for females than males. The reason for this can be attributed to the presence of more females than males in the study, leading to the skewed class distribution due to the presence of more behaviours of risk windows for females than males, enhancing the method’s learning and generalization capabilities for females.